# Verifying Equilibria in Finite-Horizon Probabilistic Concurrent Game Systems

### v4

@@ 4.2.2. Game Construction
In this section, we provide details on how to construct the $1$-bounded probabilistic concurrent game (i.e. a turn-based game) $\mathbb{G}=\langle V,v_{0},\Omega,A=\{A_{1},A_{2},\ldots A_{n+1}\},P,\delta,\mathbb{L},G=\{G_{1},G_{2},\ldots G_{n+1}\},F\rangle$
@@
      The set $V_{1}$ is $[n]\times\{\emptyset,\vee\}$, where $[n]$ denote $\{1,2\ldots n\}$. These states keep track of the position of the head and whether the current state of the machine is existential or not. Recall that when $M$ is an existential state over cell $i$, Agent $i$ needs to first be informed about which successor Agent $n+1$ wants before outputting the transition information.
@@
      The set $V_{2}$ is $[n]\times\Gamma\times\{\rightarrow,\leftarrow\}\times R$ These states correspond to the information made by a transition – the last position in $[n]$ of the head, the character in $\Gamma$ written by the head, the direction $\rightarrow$ or $\leftarrow$ the head is moving in, and the new state in $R$ of the machine. When a state $v\in V_{2}$ is visited, this symbolizes that a transition has been made in $M$. The state $v$ then contains all the necessary information for each agent in the game to accurately track the ID of $M$ through the state spaces of the individual $\pi_{i}$ transducers.
@@
      The set $V_{3}$ is $[n]\times\{\alpha,\beta\}$. These states are used to record the existential choices of Agent $n+1$. If the game was previously in a state $\langle i,\vee\rangle\in V_{1}$, then Agent $n+1$ may specify which existential successor ($\alpha$ or $\beta$) they would like to see. Once this is specified, the game transitions to either $\langle i,\alpha\rangle$ or $\langle i,\beta\rangle\in V_{3}$. This represents the aforementioned intermediate step used to deal with existential states used to relay the choices of Agent $n+1$ to the other agents.
@@
    The playing function $P$ is specified in the following way
@@
    The transition function $\delta$ is given by the following
@@
      From a state $v=\langle i,\gamma,d,r\rangle$ in $[n]\times\Gamma\times\{\rightarrow,\leftarrow\}\times R$ in $V_{2}$ the transition from $v$ is given by one of three possible cases
@@
          If the state $r$ is existential, the transition moves to $\langle i^{\prime},\vee\rangle$
@@ 4.2.3. Input Strategy Profile Construction
As mentioned before, Agent $n+1$’s assigned strategy $\pi_{n+1}$ instructs him to choose the $\alpha$ action at the ${\sf start}$ state. This gives him a probability of $1-(0.25)^{(F-1)}$ of reaching the ${\sf goal}$ state, as the only way to not reach the ${\sf goal}$ state is to loop at least $F-1$ times at the ${\sf base}$ state. The transducer implementing this strategy is of constant size.
@@
[…]
@@
    The input alphabet $V$ is the state set $V$ of the game $\mathbb{G}$ constructed in Section 4.2.2
@@
    The transition function $\gamma^{i}$ is based on component-wise updates. Let $g$ be an element of $\Gamma$, $r$ an element of $\{R\cup\emptyset\}$, and $s$ an element of $\{\alpha,\beta,\emptyset\}$. Assume the transducer $\pi_{i}$ is in state $\langle g,r,s\rangle$. We now describe how the transition function $\gamma^{i}$ behaves by describing how each of the three components transitions upon reading a state $v$.
@@
      If the input is $v=\langle i,g^{\prime},d,r^{\prime}\rangle\in V_{2}$, for $i\in[n]$, $g^{\prime}\in\Gamma$, $d^{\prime}\in\{\rightarrow,\leftarrow\}$, and $r^{\prime}\in R$. This corresponds to a new transition in the machine $M$ being observed, potentially updating the elements $g$ and $r$ in the state space of $\pi_{i}$. The state $v=\langle i,g^{\prime},d,r^{\prime}\rangle$ represents the following information about the transition in the machine $M$: the head is in cell $i$, where it writes the character $g^{\prime}$ onto the tape, transitions to the new state $r^{\prime}$, moving the head in direction $d$. The agents then use this information to update the state of their individual strategies $\pi_{i}$. Specifically, as $\pi_{i}$ is in state $\langle g,r,s\rangle$
@@
        If $v=\langle j,g^{\prime},d,r\rangle$ for a cell $j$ such that either $j=i-1$ and $d=\rightarrow$ or $j=i+1$ and $d=\leftarrow$, then the head is moving to cell $i$ in state $r^{\prime}$. So the middle component of $\langle g,r,s\rangle$ is then updated to $r^{\prime}$ to account for the new head position and state in the machine $M$’s ID.
@@
      If $\kappa(r)=\wedge$, then $r$ is a universal state and the output at state $\langle g,r,s\rangle$ is a $/50$ 50 distribution over the actions that correspond to the transitions for $r_{\alpha}$ and $r_{\beta}$, where the actions that correspond to $r_{\alpha}$ and $r_{\beta}$ are analogous to the deterministic setting above. Note that for both universal and deterministic states, we define the output function at all values of $s$ for completeness. In the actual run of the transducers, however, the only element of $s$ we see at these states is $\emptyset$, as neither deterministic nor universal states register an existential successor choice from Agent $n+1$.
@@ 4.2.4. Validity of the Reduction
Suppose that $M$ accepts the empty tape. Then, it admits an accepting computation tree $T$. By Observation 9, it also accepts a computation tree $T^{\prime}$ in which no branch contains a repeated ID. By following the strategy outlined by $T^{\prime}$ (choosing the same existential transition at an ID as $T^{\prime}$), Agent $n+1$ can meet his goal with probability $1$ in less than $F-1$ steps. Since $1>1-(0.25)^{(F-1)}$, Agent $n+1$ has an incentive to deviate from $\pi_{n+1}$ and $\pi$ is not a Nash equilibrium.
@@
5. Concluding Remarks
The two focal results of this paper are Theorem 7, which establishes a PSPACE upper bound for the subgame perfect equilibrium verification problem, and Theorem 11, which establishes an EXPTIME lower bound for the Nash equilibrium verification problem. These results contradict our intuitive understanding of the two solution concepts, as the subgame perfect equilibrium is seen as a stricter refinement of the Nash equilibrium that offers conceptual benefits at the cost of being harder to reason about. Here, we see that for the verification problem specifically, the stricter conditions of the subgame perfect equilibrium make it easier to reason about than the Nash equilibrium by allowing for efficient parallelization. When we contrast this with the fact that Nash equilibria are usually computed in these systems through the same backwards induction that creates subgame perfect equilibria, these two observations paint an interesting picture of the two equilibrium concepts that we intend to investigate further. For example, if we consider infinite-horizon systems, does this relationship still hold? If so, this increases the evidence for the conjecture that the subgame perfect equilibrium is a strictly preferable solution concept in probabilistic systems.
If we consider a deviating agent $j$, we can see the fundamental difference between the subgame perfect equilibrium and the Nash equilibrium – by examining the Markov Decision Process chain $\mathbb{G}_{i}\times\pi$ (see Section 3.4), we can see that the conditions of a strategic improvement for agent $j$ are different. For the subgame perfect equilibrium, if the value at some history-relevant state can be improved, then $\pi_{j}$ can be deviated from, and $\pi$ is not a subgame perfect equilibrium. This closely corresponds to the concept of an “optimal policy” from the literature on Markov Decision Processes [Put94], which is a memoryless strategy that maximizes the value of all states simultaneously. However, for the Nash equilibrium, the agent is only incentivized to deviate if the value of the initial state specifically can be improved, which does not directly correspond to notions of optimality in the Markov Decision Process literature. This observation paints an interesting picture of the relationship between these equilibrium concepts and other strategy-related solution concepts.

### v5

@@ 4.2.2. Game Construction
In this section, we provide details on how to construct the $1$-bounded probabilistic concurrent game (i.e. a turn-based game) $\mathbb{G}=\langle V,v_{0},\Omega,A=\{A_{1},A_{2},\ldots A_{n+1}\},P,\delta,\mathbb{L},G=\{G_{1},G_{2},\ldots G_{n+1}\},F\rangle$ .
@@
      The set $V_{1}$ is $n]\times\{\emptyset,\vee\}$, where $[n]$ denote $\{1,2\ldots n\}$. These states keep track of the position of the head and whether the current state of the machine is existential or not. Recall that when $M$ is an existential state over cell $i$, Agent $i$ needs to first be informed about which successor Agent $n+1$ wants before outputting the transition information.
@@
      The set $V_{2}$ is $n]\times\Gamma\times\{\rightarrow,\leftarrow\}\times R$ . These states correspond to the information made by a transition – the last position in $[n]$ of the head, the character in $\Gamma$ written by the head, the direction $\rightarrow$ or $\leftarrow$ the head is moving in, and the new state in $R$ of the machine. When a state $v\in V_{2}$ is visited, this symbolizes that a transition has been made in $M$. The state $v$ then contains all the necessary information for each agent in the game to accurately track the ID of $M$ through the state spaces of the individual $\pi_{i}$ transducers.
@@
      The set $V_{3}$ is $n]\times\{\alpha,\beta\}$. These states are used to record the existential choices of Agent $n+1$. If the game was previously in a state $\langle i,\vee\rangle\in V_{1}$, then Agent $n+1$ may specify which existential successor ($\alpha$ or $\beta$) they would like to see. Once this is specified, the game transitions to either $\langle i,\alpha\rangle$ or $\langle i,\beta\rangle\in V_{3}$. This represents the aforementioned intermediate step used to deal with existential states used to relay the choices of Agent $n+1$ to the other agents.
@@
    The playing function $P$ is specified in the following way:
@@
    The transition function $\delta$ is given by the following:
@@
      From a state $v=\langle i,\gamma,d,r\rangle$ in $n]\times\Gamma\times\{\rightarrow,\leftarrow\}\times R$ in $V_{2}$ the transition from $v$ is given by one of three possible cases:
@@
          If the state $r$ is existential, the transition moves to $\langle i^{\prime},\vee\rangle$.
@@ 4.2.3. Input Strategy Profile Construction
As mentioned before, Agent $n+1$’s assigned strategy $\pi_{n+1}$ instructs him to choose the $\alpha$ action at the $\sf start$} state. This gives him a probability of $1-(0.25)^{(F-1)}$ of reaching the $\sf goal$} state, as the only way to not reach the $\sf goal$} state is to loop at least F1F-1 times at the $\sf base$} state. The transducer implementing this strategy is of constant size.
@@
    The initial state of each $\pi_{i}$ corresponds to the initial configuration of the tape. Since we assumed the head started in the leftmost cell, the initial state of $\pi_{1}$ is $\langle\emptyset,r_{0},\emptyset\rangle$, corresponding to the head starting in this cell in state $r_{0}$. For the other n1n-1 agents, the initial state of $\pi_{i}$ is $\langle\emptyset,\emptyset,\emptyset\rangle$, signifying an empty cell without the head.
@@
    The input alphabet $V$ is the state set $V$ of the game $\mathbb{G}$ constructed in Section 4.2.2.
@@
    The transition function $\gamma^{i}$ is based on component-wise updates. Let $g$ be an element of $\Gamma$, $r$ an element of $\{R\cup\emptyset\}$, and $s$ an element of $\{\alpha,\beta,\emptyset\}$. Assume the transducer $\pi_{i}$ is in state $\langle g,r,s\rangle$. We now describe how the transition function $\gamma^{i}$ behaves by describing how each of the three components transitions upon reading a state $v$:
@@
      If the input is $v=\langle i,g^{\prime},d,r^{\prime}\rangle\in V_{2}$, for $i\in[n]$, $g^{\prime}\in\Gamma$, $d^{\prime}\in\{\rightarrow,\leftarrow\}$, and $r^{\prime}\in R$. This corresponds to a new transition in the machine $M$ being observed, potentially updating the elements $g$ and $r$ in the state space of $\pi_{i}$. The state $v=\langle i,g^{\prime},d,r^{\prime}\rangle$ represents the following information about the transition in the machine $M$: the head is in cell $i$, where it writes the character $g^{\prime}$ onto the tape, transitions to the new state $r^{\prime}$, moving the head in direction $d$. The agents then use this information to update the state of their individual strategies $\pi_{i}$. Specifically, as $\pi_{i}$ is in state $\langle g,r,s\rangle$:
@@
        If $v=\langle j,g^{\prime},d,r\rangle$ for a cell $j$ such that either j=i1j=i-1 and $d=\rightarrow$ or $j=i+1$ and $d=\leftarrow$, then the head is moving to cell $i$ in state $r^{\prime}$. So the middle component of $\langle g,r,s\rangle$ is then updated to $r^{\prime}$ to account for the new head position and state in the machine $M$’s ID.
@@
      If $\kappa(r)=\wedge$, then $r$ is a universal state and the output at state $\langle g,r,s\rangle$ is a $50/50$ distribution over the actions that correspond to the transitions for $r_{\alpha}$ and $r_{\beta}$, where the actions that correspond to $r_{\alpha}$ and $r_{\beta}$ are analogous to the deterministic setting above. Note that for both universal and deterministic states, we define the output function at all values of $s$ for completeness. In the actual run of the transducers, however, the only element of $s$ we see at these states is $\emptyset$, as neither deterministic nor universal states register an existential successor choice from Agent $n+1$.
@@ 4.2.4. Validity of the Reduction
Suppose that $M$ accepts the empty tape. Then, it admits an accepting computation tree $T$. By Remark 9, it also accepts a computation tree $T^{\prime}$ in which no branch contains a repeated ID. By following the strategy outlined by $T^{\prime}$ (choosing the same existential transition at an ID as $T^{\prime}$), Agent $n+1$ can meet his goal with probability $1$ in less than F1F-1 steps. Since $1>1-(0.25)^{(F-1)}$, Agent $n+1$ has an incentive to deviate from $\pi_{n+1}$ and $\pi$ is not a Nash equilibrium.
@@
5. Concluding Remarks and Future Work
The two focal results of this paper are Theorem 7, which establishes a PSPACE upper bound for the subgame-perfect-equilibrium verification problem, and Theorem 11, which establishes an EXPTIME lower bound for the Nash-equilibrium verification problem. These results contradict our intuitive understanding of the two solution concepts, as the subgame perfect equilibrium is seen as a stricter refinement of the Nash equilibrium that offers conceptual benefits at the cost of being harder to reason about. Here, we see that for the verification problem specifically, the stricter conditions of the subgame perfect equilibrium make it easier to reason about than the Nash equilibrium by allowing for efficient parallelization. When we contrast this with the fact that Nash equilibria are usually computed in these systems through the same backwards induction that creates subgame perfect equilibria, these two observations paint an interesting picture of the two equilibrium concepts that we intend to investigate in future work. For example, if we consider infinite-horizon systems, does this relationship still hold? If so, this increases the evidence for the conjecture that the subgame perfect equilibrium is a strictly preferable solution concept in probabilistic systems.
If we consider a deviating agent $j$, we can see the fundamental difference between subgame perfect equilibrium and Nash equilibrium – by examining the Markov Decision Process chain $\mathbb{G}_{i}\times\pi$ (see Section 3.4), we can see that the conditions of a strategic improvement for agent $j$ are different. For subgame perfect equilibrium, if the value at some history-relevant state can be improved, then $\pi_{j}$ can be deviated from, and $\pi$ is not a subgame perfect equilibrium. This closely corresponds to the concept of an “optimal policy” from the literature on Markov Decision Processes [Put94], which is a memoryless strategy that maximizes the value of all states simultaneously. For Nash equilibrium, however, the agent is only incentivized to deviate if the value of the initial state specifically can be improved, which does not directly correspond to notions of optimality in the Markov-Decision-Process literature. This observation paints an interesting picture of the relationship between these equilibrium concepts and other strategy-related solution concepts.